OpenStudy (anonymous):

State a rule that could be use to approximate the amout of time t needed to triple the amount of money in a savings account paying r percent interest compounded continuously.

OpenStudy (anonymous):

End result=Pe^rt We want the end result to be three time as much as beginning P. 3=1e^rt 3=e^rt

OpenStudy (anonymous):

Now we have to solve for t